OpenStudy (anonymous):

How do you write perfect square trinomials? I seem kind of incapable of understanding this concept, so could anyone explain it maybe in very simple terms, preferably with examples? Thanks.

OpenStudy (isaidavila):

ax^2+bx+C this is a perfect square trinomial bjt it can be written as for example (x+1)=x^2+2x+1

OpenStudy (isaidavila):

(x+1)^2=x^2+2x+1

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

@isaidavila Thanks for your time -- I think that helped to clarify the concept for me.

OpenStudy (isaidavila):

No problem @ireallydontlikemath

OpenStudy (welshfella):

x^2 - 8x + 16 is another this factors to (x - 4)^2